# ReflecSched: Solving Dynamic Flexible Job-Shop Scheduling via LLM-Powered Hierarchical Reflection

### v2

@@ 6.1 Experimental Setup
Benchmark Datasets. Our evaluation is performed on two benchmark suites developed for this study: GEN-Bench, for general-purpose evaluation, and PDR-Bench, a diagnostic benchmark to assess heuristic utilization. Both suites contain instances at two scales: Normal and Small. The detailed generation and curation methodology for these datasets is described in D.
@@
    $\text{RPD}(\%)=\frac{C_{\max}^{\text{method}}-C_{\max}^{\text{best}}}{C_{\max}^{\text{best}}}\times 100$ (22)
    where $C_{\max}^{\text{method}}$ is the makespan from the evaluated method and $C_{\max}^{\text{best}}$ is the best makespan found for that instance across all methods and runs in this study. Lower RPD is better.
@@
To assess statistical significance, we use the non-parametric Wilcoxon signed-rank test [Janez2006Statistical]. We evaluate a range of LLMs, including GPT-4o [openai2024gpt4ocard], DeepSeek-V3 [deepseekai2025deepseekv3technicalreport], and the Qwen3 series (8B, 14B, 32B) [yang2025qwen3technicalreport]. All models are prompted in a zero-shot manner with Chain-of-Thought reasoning [wei2022chain].
Evaluation Protocol. For initial diagnostic analyses aimed at capturing intrinsic model behaviors, we employ a multi-sample voting scheme: actions are decided by a majority vote from 5 independent generations at a high temperature (T=0.8T=0.8[wang2022self]. For the final comparative experiments against the LLM-Direct baseline, all LLM inferences use a low temperature ($T=0.2$) for efficiency and stability. To mitigate stochasticity, we report the average results across 3 independent runs.
@@
Table 3: Comparison of our ReflecSched framework against baseline methods on Normal and Small datasets. Performance is evaluated using Average Relative Percentage Deviation (RPD, %) and Average Rank. Lower values are better for both metrics.
Category Method / Model Normal Dataset Small Dataset Overall Average
Avg RPD (%) Avg Rank Avg RPD (%) Avg Rank Avg RPD (%) Avg Rank
Baseline Methods NHS 8.27 5.70 13.42 6.10 10.85 5.90
CWH 7.84 5.35 13.86 4.75 10.85 5.05
DDQN-JS 13.10 6.40 4.91 3.30 9.01 4.85
ReflecSched (Ours) Qwen3-8b 3.38 3.50 10.01 4.15 6.69 3.83
Qwen3-14b 3.75 3.70 8.33 3.25 6.04 3.48
Qwen3-32b 3.95 3.75 8.29 2.60 6.12 3.18
GPT-4o 4.08 3.70 10.24 3.55 7.16 3.63
Deepseek-v3 4.86 3.75 9.89 3.35 7.38 3.55
@@ 6.2.2 Comparison with Baseline Methods
To benchmark our ReflecSched framework, we selected and adapted three advanced algorithms representing a spectrum of established paradigms. The novelty of our problem’s specific dynamic event types necessitated this adaptation, as no directly comparable open-source implementations exist. Our comparative analysis, therefore, is not intended to challenge the original results of these baselines in their native domains, but to rigorously demonstrate the practical advantages of our proposed LLM-driven approach in this specific, complex scheduling context.
The first, CWH (Composite Weighted Heuristic), represents a sophisticated heuristic approach. Unlike the single-criterion dispatching rules evaluated previously in our study like Table 2 and Figure 6, CWH calculates job priority by integrating multiple system features through a static, linear weighted sum, providing a strong non-learning baseline. Building on this, NHS (Neural Heuristic Scorer) advances the heuristic concept by replacing the static linear weights of CWH with a neural network, allowing it to learn a non-linear scoring function and adapt its policy based on training data. Moving into the domain of DRL, we included DDQN-JS, a single-agent DRL algorithm that learns an end-to-end scheduling policy directly from environmental interactions using a Double Deep Q-Network [liu2022deep].
The performance of ReflecSched, instantiated with five different LLMs, was compared against these baselines. The results, summarized in Table 3, unequivocally demonstrate the superior performance of our framework. Our best-performing variants, ReflecSched+Qwen3-14b and ReflecSched+Qwen3-32b, achieved the best overall average RPD and average rank, respectively. This represents a substantial performance leap over the strongest baseline, DDQN-JS, which only attained an overall average RPD of 9.01%. This underscores the significant advantage provided by our framework’s planning and reflection capabilities.
On the Normal Dataset, the performance gap is most pronounced. All ReflecSched instances achieve an RPD of 4.86% or lower, whereas the best-performing baseline on this dataset CWH could only reach 7.84%. The results on the Small Dataset reveal a more nuanced picture; while DDQN-JS is competitive in terms of RPD, our ReflecSched+Qwen3-32b variant achieves the absolute best average rank across all methods and datasets, highlighting its ability to generate consistently high-quality solutions.
@@ 6.3.2 Impact of Experience Quality and Evidence Selection
Evidence Selection Strategy. We investigated four strategies for selecting historical decisions as evidence: our Full approach; Top-K Best, using only successful examples; Top-K Worst, using only failure cases; and Quartile, which samples from different performance tiers. The results on the right of Figure 8 show that none of the alternatives match the comprehensive guidance of the Full experience. Interestingly, on the Normal Dataset, the Top-K Worst strategy outperforms Top-K Best, suggesting that learning from failures can be more valuable than learning exclusively from successes. This establishes that a comprehensive evidence base provides the most robust learning signal.
@@ 6.3.3 Parameter Sensitivity Analysis under Budget Constraint
6.4 Token Consumption Analysis
A critical consideration for any LLM-based framework is its computational cost, measured by token consumption. We measured the average total tokens (input + output) consumed per instance for ReflecSched and the LLM-Direct baseline on the GEN-Bench dataset.

### v3

@@ 6.1 Experimental Setup
Benchmark Datasets. Our evaluation is performed on two benchmark suites developed for this study: GEN-Bench, for general-purpose evaluation, and PDR-Bench, a diagnostic benchmark to assess heuristic utilization. Both suites contain instances at two scales: Normal and Small. Specifically, GEN-Bench consists of 20 instances for the Normal scale and 18 instances for the Small scale, while PDR-Bench includes 111 instances for the Normal scale and 12 instances for the Small scale. The detailed generation and curation methodology for these datasets is described in D.
To evaluate cross-benchmark generalizability, we additionally report results on MK-Bench and JMS-Bench. MK-Bench is derived from the Brandimarte MK01 to MK10 instances [10], and JMS-Bench models a semiconductor cluster-tool manufacturing scenario; detailed descriptions are provided in Sections 6.3 and 6.4, respectively.
@@
    $\color[rgb]{0,0,1}\text{RPD}(\%)=\frac{C_{\max}-C_{\max}^{\star}}{C_{\max}}\times 100$ (22)
    where $C_{\max}$ is the makespan from the evaluated method and $C_{\max}^{\star}$ is the best makespan found for that instance across all methods and runs in this study. Lower RPD is better.
@@
To assess statistical significance, we use the non-parametric Wilcoxon signed-rank test [15]. We evaluate a range of LLMs, including GPT-4o [23], DeepSeek-V3 [32], and the Qwen3 series (8B, 14B, 32B) [60]. All models are prompted in a zero-shot manner with Chain-of-Thought reasoning [56].
Evaluation Protocol. For initial diagnostic analyses aimed at capturing intrinsic model behaviors, we employ a multi-sample voting scheme: actions are decided by a majority vote from 5 independent generations at a high temperature (T=0.8T=0.8[55]. For the final comparative experiments against the LLM-Direct baseline, all LLM inferences use a low temperature ($T=0.2$) for efficiency and stability.
To ensure the reliability of our findings and mitigate the stochastic nature of LLM generation, we conduct three independent runs for each problem instance and use the resulting mean values for our primary analysis. The statistical significance of the performance improvements is assessed using the non-parametric Wilcoxon signed-rank test. It is important to note that the sample size for this statistical test is defined by the total number of problem instances in the benchmark suites instead of the number of independent runs per instance. This methodology ensures that the statistical claims are based on a diverse population of scheduling scenarios, while the repeated runs for each case provide a more stable and accurate estimate of the expected performance for each individual instance.
@@
Table 3: Comprehensive performance evaluation of ReflecSched variants against baseline algorithms across diverse benchmarks. Performance is quantified via Average Relative Percentage Deviation (RPD, %) and Average Rank, where lower values signify optimal outcomes. Bold values denote the best performance in each row.
(a) General Performance Evaluation on GEN-Bench
Metric Baseline Methods ReflecSched
GP HMPSAC IDDQN DAN PPO-OC Q3-8b Q3-14b Q3-32b DS-v3 DS-v3.2 GPT-4o GPT-5n
Avg. RPD (%) 54.85 11.00 10.45 10.74 13.47 6.94 6.09 6.87 7.49 7.74 7.14 8.03
Avg. Rank 11.42 6.68 6.58 6.47 8.05 5.34 4.39 5.45 5.58 5.21 5.34 5.32
(b) Cross-Benchmark Stability on MK-Bench and JMS-Bench
Dataset Metric Baseline Methods ReflecSched
GP HMPSAC IDDQN DAN PPO-OC Q3-8b Q3-14b Q3-32b DS-v3.2 GPT-5n
MK-Bench Avg. RPD (%) 55.80 14.32 11.85 17.81 15.91 10.30 14.30 13.83 6.83 13.19
Avg. Rank 9.00 5.40 4.70 6.60 5.20 4.00 5.70 5.20 3.90 4.90
JMS-Bench Avg. RPD (%) 46.80 9.72 8.01 15.56 10.94 6.47 6.18 7.74 8.50 10.16
Avg. Rank 9.70 5.00 4.70 7.10 5.70 4.20 4.00 4.10 4.40 5.30
Note: Q3 denotes the Qwen3 series [60], DS denotes the DeepSeek series [32, 14], and GPT-5n refers to GPT-5 nano [49]. RPD values are expressed as percentages to highlight subtle performance variations between competitive models.
@@ 6.2 Comparative Performance Analysis
In this section, we conduct a detailed quantitative analysis of ReflecSched using the GEN-Bench dataset to evaluate its general performance across varying problem scales.
@@ 6.2.2 Comparison with Baseline Methods
To benchmark our ReflecSched framework, we selected and adapted five recent and competitive baselines spanning evolutionary and reinforcement learning approaches to scheduling. The baselines are taken from recent peer reviewed literature and are chosen to cover diverse modeling and optimization styles. Specifically, we include GP, an evolutionary method that uses genetic programming to evolve priority functions over a predefined feature set, reducing reliance on manually designed dispatching rules [38]. We also include DAN, an attention based model that jointly attends to operations and machines to learn instance specific representations for scheduling [54].
For benchmarking, we evaluate three distinct reinforcement learning baselines with diverse architectural characteristics. IDDQN is a Double DQN variant optimized for dispatching rule selection under stochastic machine breakdowns, employing a dueling network architecture and a modified experience replay scheme [57]. PPO-OC implements a Proximal Policy Optimization policy that maps production states to dispatching rules through a discrete action space [65]. HMPSAC adopts a hierarchical multi-policy Soft Actor-Critic framework, where a high-level controller determines intermediate objective signals and low-level policies execute specific dispatching rules [16]. Since our problem formulation involves dynamic event types that deviate from the original assumptions of these algorithms, we adapted their state representations and action interfaces to ensure compatibility with our event-driven model. The comprehensive training protocols and hyperparameter configurations for all evaluated methods are detailed in Table H.1. To ensure empirical fairness and a rigorous comparison, we strictly adhered to the canonical configurations and optimal settings recommended in the original literature for each baseline algorithm.
The performance of ReflecSched, instantiated with seven different LLM backends, was compared against these baselines. The results, summarized in Table 3, unequivocally demonstrate the superior performance of our framework. To facilitate a more granular assessment of per-instance performance and variability, we provide a comprehensive record of raw makespan results in Appendix G. This supplement includes individual data for the LLM variants, specifically Qwen3-8B, Qwen3-14B, Qwen3-32B [60], DeepSeek-V3 [32], DeepSeek-V3.2 [14], GPT-4o [23], and GPT-5 Nano [49], alongside the five specialized comparative algorithms.
As detailed in Table 3(a), all ReflecSched configurations demonstrate a substantial performance advantage over the traditional reinforcement learning and heuristic baselines on the GEN-Bench dataset. While the most competitive baseline method IDDQN achieves an average RPD of 10.45 percent, every variant of the ReflecSched framework maintains a significantly lower deviation with values ranging from 6.09 percent to 8.03 percent. Notably, the configuration utilizing the Qwen3 14b model delivers the most superior outcomes by attaining the minimum average RPD of 6.09 percent and the leading average rank of 4.39. Other specialized reinforcement learning approaches such as HMPSAC and DAN exhibit higher deviations of 11.00 percent and 10.74 percent respectively, which further underscores the efficacy of the hierarchical reflection mechanism in optimizing complex scheduling trajectories. The consistency of the proposed framework is further validated by the rank metrics where the ReflecSched variants occupy the top seven positions while the GP heuristic trails significantly with an average rank of 11.42. These results indicate that ReflecSched possesses a robust capacity for high-quality solution generation that exceeds the capabilities of models relying on direct generation or conventional training paradigms.
@@
6.3 Evaluation on Established Benchmarks: MK-Bench
To validate the cross-configuration generalizability of ReflecSched, the framework is benchmarked against MK-Bench, a suite derived from the established Brandimarte MK01–MK10 Flexible Job-Shop Scheduling instances [10]. The integration of these recognized benchmarks ensures consistency with standardized scheduling research. Given that the original Brandimarte instances are inherently static, we implement a systematic transformation protocol to extend them into dynamic environments. This adaptation introduces stochastic event arrivals while preserving the intrinsic shop topology, specifically the job and machine counts, as well as the nominal processing durations.
The adaptation protocol is structured into three sequential phases. First, the scheduling horizon for each instance is determined by deriving a theoretical workload lower bound, defined as the maximum of the longest individual job processing time and the average machine-wise workload. This bound is subsequently scaled by a factor of 1.2 to define the simulation horizon, thereby providing sufficient temporal slack for dynamic adjustments. Second, the three primary bottleneck machines are identified based on their average workload share, establishing a systematic basis for the introduction of targeted stochastic disruptions. Third, four categories of dynamic events are injected using a deterministic random seed to ensure experimental reproducibility. In these scenarios, 60% of the jobs are initialized at $t=0$ to simulate initial work-in-progress (WIP), while subsequent tasks follow an exponential arrival distribution restricted to the first half of the horizon. Furthermore, each resource is subject to a 50% breakdown probability, with mean time to repair sampled uniformly between 1.0 and 4.0 time units. Furthermore, the transformation incorporates stochastic job cancellations with a 30% probability and the insertion of high-priority urgent jobs, which arrive at timestamps sampled uniformly from 25% to 75% of the simulation horizon. These urgent tasks are strategically assigned to the previously identified bottleneck machines to intensify resource contention and evaluate the framework’s robustness under adversarial conditions.
As detailed in the MK-Bench results within Table 3(b), the ReflecSched framework demonstrates high stability when applied to flexible job-shop scheduling problems. While the most effective DRL baseline IDDQN achieves an average RPD of 11.85 percent, the ReflecSched variants maintain a clear performance advantage within this benchmarking suite. Specifically, the variant utilizing DeepSeek v3.2 attains the most favorable outcomes with an average RPD of 6.83 percent and a leading average rank of 3.90. This level of performance indicates that the hierarchical reflection mechanism successfully identifies fundamental coordination patterns that extend beyond the specific distributions encountered during the initial prompts. This finding validates the efficacy of the framework in addressing the intensive resource competition and machine flexibility constraints inherent in the standardized MK01 through MK10 instances. The overall results suggest that the strategic insights derived from the reflection module facilitate a robust approach to managing diverse workshop topologies without the necessity for specialized retraining or architecture modification.
6.4 Validation on Manufacturing Scenarios: JMS-Bench
To evaluate the industrial applicability of ReflecSched, we benchmark the framework on the JMS-Bench dataset, which is tailored to replicate the operational dynamics of a semiconductor cluster tool. In modern wafer fabrication, the cluster tool constitutes the primary production unit, comprising a central transfer robot integrated with five heterogeneous processing modules, designated as PM1 through PM5, as shown in Figure 2. Within this manufacturing context, each job represents a Front Opening Unified Pod (FOUP) carrying a batch of silicon wafers that must undergo a sequence of high-precision chemical and physical operations. The production flow is comprised of four distinct processing stages: Rough Etching, Finishing, Surface Grinding, and Metrology Inspection.
The scheduling complexity of this system is exacerbated by the functional overlap of chamber capabilities across these processing stages. Specifically, the Rough Etching stage utilizes $PM_{1}$ and $PM_{2}$, while the Finishing stage is shared between $PM_{2}$ and $PM_{3}$. Surface Grinding is distributed across $PM_{3}$ and $PM_{4}$, whereas Metrology Inspection is primarily handled by $PM_{5}$ with occasional overflow support from $PM_{4}$. This configuration dictates that $PM_{2}$, $PM_{3}$, and $PM_{4}$ serve as dual-functional modules that interlink consecutive processing phases. Consequently, the scheduler must manage acute resource contention, where the completion of a finishing task in $PM_{3}$ may preclude the initiation of a grinding operation for a concurrent job. Unlike traditional job shops, the central robot acts as a critical shared transport resource that must be tightly synchronized with these overlapping chamber cycles to prevent deadlocks, requiring the scheduler to incorporate look-ahead reasoning regarding module availability to sustain high throughput.
In this case study, we characterize three representative categories of dynamic disturbances inherent to semiconductor manufacturing: the insertion of high-priority “hot lots” necessitating immediate preemption of nominal production, stochastic chamber failures induced by equipment drift, and job cancellations arising from upstream yield fluctuations. These events necessitate non-myopic reasoning to re-allocate resources while preserving the integrity of the global production schedule. The Strategic Experience produced by ReflecSched is designed to encode such high level coordination patterns, for example by deciding when a shared resource such as $PM_{4}$ should be kept available for metrology to reduce queue buildup at the inspection stage during high arrival periods.
The comparative performance metrics for the semiconductor manufacturing environment are detailed in Subtable b of Table 3(b). According to the recorded data, the ReflecSched variants demonstrate significant advantages in managing the complex operational constraints of cluster tools when compared to both meta-heuristic and reinforcement learning baselines. Among the baseline algorithms, IDDQN emerges as the most competitive approach with an average RPD of 8.01 percent and an average rank of 4.70. However, the ReflecSched configuration powered by the Qwen3 14b model surpasses all baselines by achieving a minimum average RPD of 6.18 percent and an optimal average rank of 4.00. Other variants such as Qwen3 8b and Qwen3 32b also maintain high solution quality with RPD values of 6.47 percent and 7.74 percent respectively. This performance trend indicates that the hierarchical reflection mechanism effectively internalizes the intricate module to stage mappings and transport synchronization requirements inherent in semiconductor fabrication. Furthermore, the proposed framework consistently outperforms specialized reinforcement learning architectures such as HMPSAC and PPO-OC which yield higher RPD values of 9.72 percent and 10.94 percent respectively. These findings suggest that the strategic insights codified within the reflection module provide a robust foundation for mitigating resource contention even under the severe stochastic disruptions characteristic of modern wafer processing units.
@@ 6.5.2 Impact of Experience Quality and Evidence Selection
Evidence Selection Strategy. We investigated four strategies for selecting the trajectories that serve as the empirical basis for reflection. Our Full approach provides contrastive insights by pairing the best and worst outcomes from the simulation pool. In contrast, Top-K Best focuses exclusively on successful trajectories to highlight optimal patterns, while Top-K Worst utilizes only the poorest-performing sequences to identify critical anti-patterns. The Quartile strategy employs a stratified sampling method to gather evidence from across the top, middle, and bottom performance tiers of the simulation results. The results on the right of Figure 9 show that none of the alternatives match the comprehensive guidance of the Full experience. Interestingly, on the Normal Dataset, the Top-K Worst strategy outperforms Top-K Best, suggesting that learning from failures to identify actions that should be avoided can be more valuable than learning exclusively from successes. This establishes that a comprehensive evidence base utilizing contrastive extremal outcomes provides the most robust learning signal.
@@ 6.5.3 Parameter Sensitivity Analysis under Budget Constraint
Figure 11: Comparative Gantt chart analysis of the semiconductor cluster tool scheduling instance. The top chart illustrates the schedule generated by the myopic baseline which achieves a makespan of 40.0. The bottom chart depicts the schedule produced by the ReflecSched framework yielding a superior makespan of 36.67. Red cross symbols denote the occurrence of dynamic machine breakdowns on modules PM3 and PM5. This comparison demonstrates how the proposed hierarchical reflection mechanism identifies global strategic trade-offs to mitigate the impact of stochastic disruptions.
@@
6.6 Case Study: Dynamic Scheduling under Stochastic Disruptions
To provide a detailed elucidation of the decision-making mechanisms within ReflecSched, we conduct a comparative analysis of two distinct scheduling trajectories for a representative production instance. This scenario comprises five processing modules ($PM_{1}$ to $PM_{5}$) and 15 wafer batches with heterogeneous processing requirements. The primary objective is to evaluate the framework’s resilience to stochastic machine failures and its ability to circumvent sub-optimal local states typical of reactive heuristics. As evidenced by the trajectory comparison, the ReflecSched-guided policy (Case A) yields an optimized makespan of 36.67, whereas the reactive baseline (Case B) results in an extended completion time of 40.0.
The dynamic scenario in this case study is characterized by two discrete equipment malfunctions: $PM_{5}$ undergoes a failure at $t=8.78$ and is restored to operational status at $t=10.74$, followed by a secondary failure of $PM_{3}$ at $t=14.80$, with recovery at $t=17.41$. Upon the initial failure of $PM_{5}$, the Strategic Experience codified by the reflection module designates Job 3 and Job 6 as critical-path tasks for the Metrology stage. Consequently, the scheduler preempts non-critical tasks, specifically Job 14 on $PM_{1}$ and Job 9 on $PM_{2}$, to expedite the resumption of Job 8 and Job 5 as resources become available. By tolerating localized delays for standard batches, ReflecSched prevents downstream delay propagation for Job 16 and Job 1, facilitating the conclusion of the final metrology operation for Job 1 at $t=36.67$
In contrast, Case B exemplifies the inherent limitations of a reactive, greedy baseline. While this approach maximizes chamber utilization during the initial 10 time units, it lacks sufficient look-ahead capability to mitigate long-term resource contention. When the identical disruptions affect $PM_{3}$ and $PM_{5}$, the baseline model fails to reconfigure the global schedule optimally. Specifically, the myopic assignment of Job 13 and Job 11 to $PM_{3}$ early in the horizon induces severe congestion at the finishing stage. This inflexible allocation logic necessitates that Job 17 remains idle until the final phases of the simulation horizon before accessing its terminal processing chamber. Consequently, the absence of strategic coordination causes Job 17 to become the primary system bottleneck, yielding an extended makespan of 40.0. This comparison provides empirical evidence that the integration of hierarchical reflection enables the scheduler to anticipate downstream bottlenecks and sustain high throughput in the presence of intensive stochastic disturbances.
6.7 Efficiency and Runtime Analysis
To evaluate the practical utility of ReflecSched, we conduct a multi-dimensional efficiency analysis encompassing wall-clock time, hardware overhead, and token consumption. All experiments were executed on a workstation featuring an Intel Xeon Platinum 8468V CPU and an NVIDIA H100 GPU (80 GB VRAM). LLM-based models were deployed via vLLM to ensure high-throughput inference [27].
In dynamic manufacturing environments, the selection of a scheduling framework requires a careful trade-off between deployment readiness and per-instance execution speed. While DRL models exhibit extremely low inference latency once weights are fully trained, their high initial training overhead serves as a significant bottleneck when factory configurations or optimization objectives shift. Conversely, ReflecSched utilizes a zero-shot architecture that enables immediate deployment without any prior training phase.
We analyze this trade-off using Cumulative Wall-clock Time as defined by the following formulation:
$T_{cum}(N)=T_{train}+\sum_{i=1}^{N}T_{infer,i}$ (23)
where $T_{train}$ represents the offline training time and $T_{infer,i}$ denotes the wall-clock time required for each scheduling instance. As illustrated in Figure 12, ReflecSched maintains a clear efficiency advantage in scenarios involving limited batch sizes or frequent environment re-configurations. Specifically, for the DeepSeek-V3.2 and GPT-5-nano backends, the cumulative time remains lower than that of the DRL baseline until the number of instances reaches a break-even point of approximately 74 and 50 respectively.
However, we must acknowledge that due to the inherent computational complexity of large language model inference and iterative reflection, the marginal time cost per instance for ReflecSched is higher than that of specialized DRL models. In static environments characterized by massive, repetitive instance batches where $N$ significantly exceeds the break-even point, the initial training investment of DRL is effectively amortized, which allows its fast inference capability to become a dominant advantage. Therefore, ReflecSched is most suitably positioned as a robust solution for high-variability manufacturing systems where the cost of repeated model retraining would be prohibitive.
Figure 12: Break-even analysis of cumulative wall-clock time. Intercepts at $N=0$ represent initial training overhead. Solid and dashed lines denote recorded and extrapolated runtimes, respectively, with markers highlighting break-even points against the HMPSAC baseline.